if A = cos(xy)i + ( 3xy + 2x ^2 )j - (3x + 2y )k find d^2A/dy^2 and d^2A/dxdy


1
Expert's answer
2021-03-31T01:53:51-0400

"\frac{dA}{dy}=-xsin(xy)i+3xj-2k"


"\frac{d^2A}{dy^2}=-x^2cos(xy)i"


"\frac{d^2A}{dxdy}=(-sin(xy)-xycos(xy))i+3j"